# A New Era in Computational Pathology: A Survey on Foundation and Vision-Language Models

### v1

@@ III-B Pre-training Workflow and Strategies
In this single diagram, both pre-training strategy for VFM and VLFM is shown. For VFMs, a vision module with some initial weight is leveraged in pre-training. The term “vision module” is used to generally represent a wide range of vision architectures and also modified versions of these architectures with additional layers. The most common architecture is variants of vision image transformers (ViTs) [78] which includes ViT-S, ViT-B, ViT-L and ViT-H. Though there are specialized architectures proposed in BEPH [79] and Prov-GigaPath [26] which uses BEiTv2 and GigaPath architecture respectively. Note that, some models initialize the architectures with ImageNet [80] weights to get an initial vision representation space which can be transformed through pre-training. In Table III, a summary of pre-training strategies in different phases is outlined.
@@
Contrastive Learning MoCov3 [83]
@@
Vision-language pre-training CoCa [86], CLIP [29]
@@
Prov-GigaPath [26] MIM and Self-Distillation MAE, DINOv2 GigaPath [26] (constructed with ViT and LongNet [95]) Providence Health System (PHS) Slides: 171,189 Tiles: 1.3 billion H&E IHC 30,000 patients 28 cancer centers 31 major tissue types Vision Language
CONCH [24] MIM + Self-Distillation iBOT ViT-B backbone in image encoder In-house dataset Slides: 21,442 Tiles: 1.3 billion Self-Distillaton 350 cancer subtypes Vision Language
The term “unimodal” is used to signify the pre-training phase utilizing a single modality out of vision and language. This is inherently different from vision-language pre-training strategies which involve both vision and language modalities to learn a joint vision-language representation space. In the unimodal vision pre-training phase there are three strategies commonly used in CPath which are self-distillation, contrastive learning and masked image modeling (MIM) approach. Each approach has its own advantages and disadvantages; however, the most popular approach is self-distillation with no labels (specifically DINOv2) which has been used in all the recent FMs [23, 64, 90, 88] as the SSL scheme. The work done in Lunit [96] goes through and benchmarks DINO, MoCov2, SwaV [97], Barlow Twins [98], but concludes there is no clear best SSL scheme. Lai [99] et al. is another work that performs a thorough evaluation of SSL schemes in CPath. .
Figure 10: Visualization of the self-distillation approach of unimodal vision pre-training scheme.
@@ III-B1 Unimodal Vision Pre-training
PathChat [25] UNI [90] as vision encoder backbone Pre-trained Llama 2 [101] LLM which is a decoder-only transformer-based auto-regressive model A multimodal projector module to connect the outputs of the vision module to language module by projecting the visual tokens to the same dimension as the LLM’s embedding space for text tokens CoCa Vision-language pre-training according to CoCa framework with CONCH dataset LLaVa 1.5 [63] training approach First phase: Only the parameters of multi-modal projector is updated Second phase: Fine-tuned with instruction-following data
@@
my tier/.style=tier/.wrap pgfmath arg=level##1level(), , The self-distillation with no label approach uses a student-teacher network to learn a rich vision representation space as shown in Fig. 10. From a single image patch, two different views are generated by applying an augmentation sampled from a set of possible augmentations (color jittering, Gaussian blur, polarization, etc). The generated output of both the networks is utilized to compute a cross-entropy loss which is then used to update the parameters of the student network. The parameter of the teacher network is then updated through an exponential moving average (EMA) of the student network parameters. Among the surveyed articles Virchow2, Virchow, RudolfV, Hibou and UNI use DINOv2 as the self-distillation approach. PLUTO takes a unique approach by integrating MAE and a Fourier loss term to get a custom variation of DINOv2.
Figure 11: Visualization of the masked image modeling approach of unimodal vision pre-training scheme.
Figure 12: Visualization of the MoCo approach of unimodal vision pre-training scheme.
The second popular SSPT approach is the MIM (visualized in Fig. 11) which has variants like MAE and iBOT used in FMs. In the MAE approach, randomly selected high portions of the image are masked out and the patches that are not masked out are passed through an encoder which generates latent representations of those patches.
@@
Among the surveyed articles 3B-CPath and Prov-GigaPath use MAE but in conjunction with DINO and DINOv2 respectively. Research works utilizing iBOT approach include Phikon and CONCH.
@@
Like the self-distillation approach, MoCo also utilizes two models; one is an encoder (with query patch as input) and the second one is a special momentum encoder (with key patches as input). The embeddings generated through the encoder and the momentum encoder are used to compute a similarity score which in turn is used for contrastive loss computation. The computed loss is used in backpropagation to update the parameters of the encoder. The parameter of the momentum encoder is updated through a momentum-based update rule that utilizes the parameters of the encoder. Another innovation of MoCo was introducing a feature queue which ensures a large set of negative samples without holding the entire dataset in memory. This also ensures the negative samples do not go stale over the training period and the model sees a diverse set of negative samples to learn a better vision representation. Among the surveyed articles PathoDuet uses MoCov3, UNI uses MoCov3 for comparison and CTransPath performs a unique variation of MoCov3 called semantically relevant contrastive learning (SRCL).
GPFM [22] is distinct from other approaches as it includes a novel expert knowledge distillation in addition to MIM and self-distillation by utilizing existing FMs UNI, Phikon and CONCH.
III-B2 Vision-Language Pre-training
After unimodal pre-training for the image module and text module, VLFMs go through vision language SSPT to align or learn a joint vision-language representation space.
@@ III-B2 Vision-Language Pre-training
Figure 13: Visualization of the CLIP approach of the vision-language pre-training scheme.
@@
Figure 14: Visualization of the CoCa approach of the vision-language pre-training scheme.
@@
One recent work Zhou et al. [^1] https://github.com/MAGIC-AI4Med/KEP [106] proposes an entirely new approach in visual-language SSPT named knowledge enhanced pre-training (KEP) by utilizing PathKT, a pathology knowledge tree curated by the same research. This is different from the CLIP and CoCa pre-training techniques as it uses a novel knowledge encoder and knowledge distillation technique.
@@ III-C Instruction-Tuning Phase
For CPath, the instruction-tuning phase provides the model with conversational ability i.e. a user can prompt the model and the model will respond according to the prompt. Visual-instruction tuning adds the ability to provide an image in addition to natural language user prompts. Among the surveyed papers in FMs, PathChat and PathAsst perform instruction-tuning with self-curated datasets PathChatInstruct and PathInstruct, respectively. Both works adopt the strategy employed by LLaVA and LLaVa-1.5 which are the pioneering work in visual instruction tuning. A summary of the instruction tuning process is provided in Table V. For both PathChat and PathAsst, the instruction tuning phase is subdivided into two phases. For both models in the first phase, the parameters of the vision encoder are frozen and the layer/module connected to the vision encoder (for PathChat it is a multimodal projector, for PathAsst it is fully connected layers) is trained. In the second phase, the model is fine-tuned with instruction-following data.
Even though instruction tuning in CPath is relatively new, some recent works like PathInsight [107] specifically focus on this aspect of FMs. Another recent work CLOVER [^2] https://github.com/JLINEkai/CLOVER [108] provides a framework for cost-effective instruction learning in pathology.
III-D Downstream Tasks
@@
  1. 1.
  2. 2.
    KNN Probing: This is yet another approach to adopt a foundation model for a specific downstream task by utilizing K-nearest neighbors algorithm.
  3. 3.
    Fine-Tuning: This is similar to linear probing as a classifier/regressor is added on top of the pre-trained model, but the major difference is the parameter of the pre-trained models is also updated during fine-tuning. Hence, it is computationally much more costly compared to linear probing. This is sometimes also referred to as the supervised training phase. Note that, most of the time annotation is only available on slide-level. Hence, if the slide-level label is utilized it is called weakly supervised training.
@@
In Table VI, a summary of downstream tasks is provided along with research works performing these tasks.
@@
    Few Shot Evaluation: In few shot evaluation, the pre-trained model sees only a few examples from the downstream task dataset.
@@
    Simple Shot Evaluation: This is a unique variation of the few shot evaluation methods which is only been explored in UNI.
@@ III-E Framework, Benchmarking and Adaptation of FMs
PanNuke [109] [40]
DigestPath [110] [40], [24]
WSSS4LUAD [111] [40],[38],[24]
SkinCancer [112] [39]
PatchCamelyon (PCam) [113] [88],[64],[23],[39]
MHIST [114] [88],[64],[23],[39]
BACH [115] [79],[90],[39]
SICAP, SICAPv2 [116] [24],[39]
Databiox [117] [39]
RenalCell [118] [39]
Osteo [119] [39]
CoNSeP [120] [23]
KIMIA Path24C [121] [40]
DHMC [122] [24],[90]
EBRAINS [123] [24],[90]
AGGC [124] [24],[90]
PANDA [125] [24],[90]
MSK-IMPACT [126] [102],[23]
SegPath [127] [90]
BRACS [128] [90]
UniToPatho [129] [90]
HunCRC [130] [90]
BreakHis [131] [79]
MSI-CRC & MSI-STAD [132, 133] [88],[64]
TIL-DET [134] [88],[64]
MIDOG [135] [23]
CAMELYON16 [136] [79],[92],[90]
CAMELYON17-WILDS [137] [90],[23]
@@
III-E1 Frameworks:
eva [^3] https://github.com/kaiko-ai/eva [71, 142] is a framework for VFMs in CPath which abstracts a lot of complexity of VFMs. In addition, it facilitates the reproducibility of VFMs for fair comparison and provides an interface to evaluate publicly available downstream datasets. Note that, this is an ongoing work and more downstream datasets and models are being added.
III-E2 Benchmarking:
Another category is benchmark analysis of FMs [67], [66],[143] in CPath. The work done in [67] analyzes different FMs and evaluates their performance on $88$ datasets. The other work [66] is a more recent work that analyzes the performance of FMs on a large and diverse data set collected from two medical centers. The FMs were benchmarked on $33$ broad downstream tasks which include disease detection, biomarker prediction and treatment outcome prediction.
III-E3 Adaptation of FMs:
Another category is the adaptation of existing FMs to carry out tasks such as low-resource fine-tuning[72] and multi-modal prompt-tuning[73]. The work carried out in [72] fine-tunes a FM with a single GPU and shows that it can outperform SOTA feature extractors. PathoTune [73] adapts a visual or pathological FM to downstream tasks using multi-modal prompts.
Other than the surveyed FMs, there are more recent works of FM in CPath like mSTAR [144]. However, as it includes RNA-Seq data in addition to pathology reports it falls outside the scope of this paper. Another foundation model we do not include is H-optimus-0444https://huggingface.co/bioptimus/H-optimus-0 [145] developed by Bioptimus as it is not associated with a publication.
@@ III-E3 Adaptation of FMs:
CLIPath [167] ResNet-50 as the vision encoder in CLIP $44$ fully-connected layer after the vision encoder Transformer as the text encoder in CLIP PatchCamelyon (PCam) MHIST Develops residual feature connection (RFC) to fine-tune CLIP with a small amount of trainable parameters and this also fuses the existing knowledge from pre-trained CLIP and the newly learned knowledge related to pathological tasks
@@
PathGen-CLIP [37] PathGen-LLaVA OpenCLIP framework is utilized to train a pathology-specific CLIP model The vision encoder of LLaVA v1.5 is replaced with the trained pathology-specific CLIP model PathGen-CLIP is combined with Vicuna LLM PathGen-1.6M (for pre-training) PatchCamelyon (Pcam) CRC-100K SICAPv2 BACH Osteo SkinCancer LC25000 Camelyon 16 and 17 Curates a large-scale dataset PathGen-1.6M with 1.6 million image-caption pairs Develops a pathology-specific large multi-modal modal with the capability to adapt to various downstream task
@@
Hu et al. [184] ResNet-50 pre-trained with BYOL [185] SSL scheme as patch encoder Anchor-based module as WSI encoder based on [186] which uses kernel attention transformer Prompt-based text encoder which utilizes self-attention structure in BERT [187] GastricADC [188] In-house gastric dataset containing 3598 WSIs and reports dataset Performs cross-modal retrieval tasks which includes image-to-image retrieval, image-to-text retrieval, text-to-image retrieval and text-to-text retrieval Proposes a histopathology language-image representation learning framework Develops a prompt-based text representation learning scheme

### v2

@@ III-B Pre-training Workflow and Strategies
In this single diagram, both pre-training strategy for VFM and VLFM is shown. For VFMs, a vision module with some initial weight is leveraged in pre-training. The term “vision module” is used to generally represent a wide range of vision architectures and also modified versions of these architectures with additional layers. The most common architecture is variants of vision image transformers (ViTs) [88] which includes ViT-S, ViT-B, ViT-L, ViT-G and ViT-H. Though there are specialized architectures proposed in BEPH [28] and Prov-GigaPath [38] which uses BEiTv2 [89] and GigaPath architecture respectively. Note that, some models initialize the architectures with ImageNet [90] weights to get an initial vision representation space which can be transformed through pre-training. In Table III, a summary of pre-training strategies in different phases is outlined.
@@
Contrastive Learning MoCov2 [93], MoCov3 [94]
@@
Vision-language pre-training CLIP [45], CoCa [97]
@@
Prov-GigaPath [38] MIM and Self-Distillation MAE, DINOv2 GigaPath [38] (constructed with ViT and LongNet [99]) Providence Health System (PHS) Slides: 171,189 Tiles: 1.3 billion H&E, IHC 30,000 patients 28 cancer centers 31 major tissue types Vision Language
CONCH [41] MIM + Self-Distillation iBOT ViT-B backbone in image encoder In-house dataset Slides: 21,442 Tiles: 1.2 million H&E, IHC, Masson’s trichrome, Congo red 350 cancer subtypes Vision Language
The term “unimodal” is used to signify the pre-training phase utilizing a single modality out of vision and language. This is inherently different from vision-language pre-training strategies which involve both vision and language modalities to learn a joint vision-language representation space. In the unimodal vision pre-training phase there are three strategies commonly used in CPath which are self-distillation, contrastive learning and masked image modeling (MIM) approach. Each approach has its own advantages and disadvantages.
Figure 10: Visualization of the self-distillation approach for unimodal vision pre-training scheme.
The work done in Lunit[34] benchmarks DINO [91], MoCov2 [93], SwaV [100], Barlow Twins [101] but concludes there is no clear best SSPT scheme. However, the authors conclusively prove that SSPT always outperforms ImageNet initialization.
@@ III-B1 Unimodal Vision Pre-training
PathChat [37] UNI [31] as vision encoder backbone Pre-trained Llama 2 [102] LLM which is a decoder-only transformer-based auto-regressive model A multimodal projector module to connect the outputs of the vision module to language module by projecting the visual tokens to the same dimension as the LLM’s embedding space for text tokens CoCa Vision-language pre-training according to CoCa framework with CONCH dataset LLaVA-1.5 [75] training approach First phase: Only the parameters of multi-modal projector is updated Second phase: Fine-tuned with instruction-following data
@@
my tier/.style=tier/.wrap pgfmath arg=level##1level(), , The self-distillation with no label approach uses a student-teacher network to learn a rich vision representation space as shown in Fig. 10. From a single image patch, two different views are generated by applying an augmentation sampled from a set of possible augmentations (color jittering, Gaussian blur, polarization, etc). The generated output of both the networks is utilized to compute a cross-entropy loss which is then used to update the parameters of the student network. The parameter of the teacher network is then updated through an exponential moving average (EMA) of the student network parameters. Among the surveyed articles Virchow2 [22], Virchow [24], RudolfV [26], Hibou [27] and UNI [31] use DINOv2 [92] as the self-distillation approach. PLUTO [29] takes a unique approach by integrating MAE and a Fourier loss term to get a custom variation of DINOv2.
Figure 11: Visualization of the masked image modeling approach for unimodal vision pre-training scheme.
The second popular SSPT approach is the MIM (visualized in Fig. 11) which has variants like MAE [95] and iBOT [96] used in FMs. In the MAE approach, randomly selected high portions of the image are masked out and the patches that are not masked out are passed through an encoder which generates latent representations of those patches.
@@
Among the surveyed articles 3B-CPath [32] and Prov-GigaPath [38] use MAE but in conjunction with DINO and DINOv2 respectively. Research works utilizing iBOT approach include Phikon [33] and CONCH [41].
@@
Like the self-distillation approach, MoCo also utilizes two models; one is an encoder (with query patch as input) and the second one is a special momentum encoder (with key patches as input). The embeddings generated through the encoder and the momentum encoder are used to compute a similarity score which in turn is used for contrastive loss computation. The computed loss is used in backpropagation to update the parameters of the encoder. The parameter of the momentum encoder is updated through a momentum-based update rule that utilizes the parameters of the encoder.
Figure 12: Visualization of the MoCo approach for unimodal vision pre-training scheme.
Another innovation of MoCo was introducing a feature queue which ensures a large set of negative samples without holding the entire dataset in memory. This also ensures the negative samples do not go stale over the training period and the model sees a diverse set of negative samples to learn a better vision representation. Among the surveyed articles PathoDuet [23] uses MoCov3, UNI [31] uses MoCov3 for comparison and CTransPath [35] performs a unique variation of MoCov3 called semantically relevant contrastive learning (SRCL). GPFM [36] is distinct from other approaches as it includes a novel expert knowledge distillation in addition to MIM and self-distillation by utilizing existing FMs UNI, Phikon and CONCH.
III-B2 Unimodal Language Pre-Training
Most VLFMs do not perform large-scale unimodal language pre-training. Instead, they rely on already pre-trained in-domain (trained with medical corpus) LLMs like BioGPT or other LLMs like Llama 2, GPT-variants, etc. One exception is CONCH which pre-trains the language module with 550,000550000550,000550 , 000 surgical pathology reports from Massachusetts General Hospital and over 400,000400000400,000400 , 000 select histopathology-relevant PubMed abstracts.
III-B3 Vision-Language Pre-training
After unimodal pre-training for the image module and text module, VLFMs might go through vision language SSPT to align or learn a joint vision-language representation space.
@@ III-B3 Vision-Language Pre-training
Figure 13: Visualization of the CLIP approach for the vision-language pre-training scheme.
@@
Figure 14: Visualization of the CoCa approach for the vision-language pre-training scheme.
@@
One recent work KEP [30] proposes an entirely new approach in visual-language SSPT named knowledge enhanced pre-training by utilizing PathKT, a pathology knowledge tree curated by the same research. This is different from the CLIP and CoCa pre-training techniques as it uses a novel knowledge encoder and knowledge distillation technique.
@@ III-C Instruction-Tuning Phase
For CPath, the instruction-tuning phase provides the model with conversational ability i.e. a user can prompt the model and the model will respond according to the prompt. Visual-instruction tuning adds the ability to provide an image in addition to natural language user prompts. Among the surveyed papers in FMs, PathChat and PathAsst perform instruction-tuning with self-curated datasets PathChatInstruct and PathInstruct, respectively. Both works adopt the strategy employed by LLaVA and LLaVA-1.5 which are the pioneering work in visual instruction tuning. A summary of the instruction tuning process is provided in Table V. For both PathChat and PathAsst, the instruction tuning phase is subdivided into two phases. For both models in the first phase, the parameters of the vision encoder are frozen and the layer/module connected to the vision encoder (for PathChat it is a multimodal projector, for PathAsst it is fully connected layers) is trained. In the second phase, the model is fine-tuned with instruction-following data. More details about visual instruction tuning and a description of research work exclusively focusing on this aspect in CPath are provided in section IV-C.
III-D Downstream Tasks and Datasets
@@
  1. 1.
  2. 2.
    KNN Probing: This is yet another approach to adopt a FM for a specific downstream task by utilizing K-nearest neighbors algorithm. In KNN probing a specific representation from learned representation space (through SSPT) is used to find its nearest neighbors.
@@
  3. 3.
    Fine-Tuning: This is similar to linear probing as a classifier/regressor is added on top of the pre-trained model, but the major difference is the parameter of the pre-trained models is also updated during fine-tuning. Hence, it is computationally much more costly compared to linear probing. This is sometimes also referred to as the supervised training phase. Note that, most of the time annotation is only available on slide-level. Hence, if the slide-level label is utilized it is called weakly supervised training. To attain the slide-level representation from the trained model (which is trained to generate patch-level embeddings), typically a aggregator model is used. A visualization of attaining the slide level representation from the patch embeddings is shown in Fig. 15.
In Table VI, a summary of downstream tasks is provided along with research works performing these tasks. In Table VII a list of downstream datasets utilized by different FMs is provided.
@@
    Few Shot Evaluation: In few shot evaluation, the pre-trained model sees only a few examples from the downstream task dataset. This is sometimes mentioned as K-shot evaluation where K is the number of data points seen by the model.
@@
    Simple Shot Evaluation: Simpleshot [106] is a unique variation of the few shot evaluation methods which is only been explored in UNI. It combines few shot learning with a nearest-neighbor classifier.
A summary visualization for these adaptation and evaluation strategies is given in Fig. 16.
@@ III-E Framework, Benchmarking and Adaptation of FMs
PanNuke [42]
DigestPath [42], [41]
WSSS4LUAD [42],[40],[41]
PatchCamelyon (PCam) [27],[26],[24]
MHIST [27],[26],[24]
BACH [28],[31]
SICAP, SICAPv2 [41]
CoNSeP [24]
KIMIA Path24C [42]
DHMC [41],[31]
EBRAINS [41],[31]
AGGC [41],[31]
PANDA [41],[31]
MSK-IMPACT [39],[24]
SegPath [31]
BRACS [31]
UniToPatho [31]
HunCRC [31]
BreakHis [28]
MSI-CRC & MSI-STAD [27],[26]
TIL-DET [27],[26]
MIDOG [24]
CAMELYON16 [28],[23],[31]
CAMELYON17-WILDS [31],[24]
@@
III-E1 Frameworks
eva [83, 107] is a framework for VFMs in CPath which abstracts a lot of complexity of VFMs. In addition, it facilitates the reproducibility of VFMs for fair comparison and provides an interface to evaluate publicly available downstream datasets.
III-E2 Benchmarking
Another category is benchmark analysis of FMs [76, 77, 79, 80, 82, 85, 87] in CPath. The work in [76] benchmarks the performance of $66$ foundation models (CTransPath, PathoDuet, PLIP, CONCH and UNI) across $55$ clinically relevant prediction tasks. The contribution of [77] is benchmarking FMs which include CTransPath, Lunit, Phikon, PLIP, UNI and Virchow for ovarian cancer bevacizumab reatment response prediction. Another work [79] analyzes the performance of FMs on a large and diverse data set collected from two medical centers. The FMs (CTransPath, UNI, Virchow and Prov-GigaPath) were benchmarked on $33$ broad downstream tasks which include disease detection, biomarker prediction and treatment outcome prediction. In [80], the authors benchmarks PathCLP [40] on various corrupted images (with $77$ corruption types) for $22$ downstream datasets. The authors of [82] investigate how layer selection for FMs (CTransPath, Phikon, Lunit + other encoders) affects the downstream task performance. The work done in [85] analyzes different FMs and evaluates their performance on $88$ datasets. In [87], the authors evaluates the performance of histopathology-specific SSL methods for $1717$ unique tissue types and $1212$ unique cancer types.
III-E3 Adaptation of FMs
Another category is the adaptation of existing FMs to carry out tasks such as low-resource fine-tuning[86] and multi-modal prompt-tuning[78]. The work carried out in [86] fine-tunes a FM with a single GPU and shows that it can outperform SOTA feature extractors. PathoTune [78] adapts a visual or pathology-specific FM to downstream tasks using multi-modal prompts. In [81] the authors proposes a task-specific visual prompting approach to tune VFMs. The work done in [84] utilizes in-context learning to learn from prompts without parameter updates. They specifically utilizes GPT-4 with vision capability and performs evaluation on $33$ downstream datasets.
Other than the surveyed FMs, there are more recent works of FM in CPath like mSTAR [108]. However, as it includes RNA-Seq data in addition to pathology reports it falls outside the scope of this paper. Another foundation model we do not include is H-optimus-0 [25] developed by Bioptimus as it is not associated with a publication.
@@ III-E3 Adaptation of FMs
CLIPath [130] ResNet-50 as the vision encoder in CLIP $44$ fully-connected layer after the vision encoder Transformer as the text encoder in CLIP PatchCamelyon (PCam) MHIST Develops residual feature connection (RFC) to fine-tune CLIP with a small amount of trainable parameters and this also fuses the existing knowledge from pre-trained CLIP and the newly learned knowledge related to pathology-specific tasks
@@
PathGen-CLIP [51] OpenCLIP framework is utilized to train a pathology-specific CLIP model The vision encoder of LLaVA v1.5 is replaced with the trained pathology-specific CLIP model PathGen-CLIP is combined with Vicuna LLM PathGen-1.6M (for pre-training) PatchCamelyon (Pcam) CRC-100K SICAPv2 BACH Osteo SkinCancer LC25000 Camelyon 16 and 17 Curates a large-scale dataset PathGen-1.6M with 1.6 million image-caption pairs Develops a pathology-specific large multi-modal modal with the capability to adapt to various downstream task
@@
Hu et al. [147] ResNet-50 pre-trained with BYOL [148] SSL scheme as patch encoder Anchor-based module as WSI encoder based on [149] which uses kernel attention transformer Prompt-based text encoder which utilizes self-attention structure in BERT [150] GastricADC In-house gastric dataset containing 3598 WSIs and reports dataset Performs cross-modal retrieval tasks which includes image-to-image retrieval, image-to-text retrieval, text-to-image retrieval and text-to-text retrieval Proposes a histopathology language-image representation learning framework Develops a prompt-based text representation learning scheme